OpenStudy (vvfvf):

Problem Page Find the slope of the line passing through the points , 6−9 and , 3−2 .

OpenStudy (foxesrcool):

screenshot plz

OpenStudy (mayankdevnani):

what do you think? any ideas?

OpenStudy (mayankdevnani):

@vvfvf you there ?

OpenStudy (foxesrcool):

No idea lol

OpenStudy (foxesrcool):

@pooja195

OpenStudy (mayankdevnani):

lol

OpenStudy (mayankdevnani):

suppose if any two points are given, (a,b) and (x,y) \[\large \bf slope=\frac{y-b}{x-a}\]

OpenStudy (mayankdevnani):

similarly,solve your question

pooja195 (pooja195):

@vvfvf Do you still need help?

OpenStudy (vvfvf):

my fault guys my mom had me do something for her but yeah

pooja195 (pooja195):

Use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \]

OpenStudy (vvfvf):

so i put 6 with 3 and -9 with -2?

pooja195 (pooja195):

yep goes in like this: \[\huge~\rm~\bf\color{blue}{ \frac{-2-(-9)}{3-6}=slope}\] two negatives make a positive so.. \[\huge~\rm~\bf\color{blue}{ \frac{-2+9}{3-6}=slope}\] and from here you could find the slope :)

OpenStudy (vvfvf):

7 - 3?

pooja195 (pooja195):

Perfect :)

OpenStudy (vvfvf):

Aye thank you

pooja195 (pooja195):

you're welcome :-)